# Overcomplete Tensor Decomposition via Koszul–Young Flattenings

### v1

@@ 1 Introduction
A tensor is a multi-way array $T\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{k}}$ with entries denoted $T_{i_{1},i_{2},\ldots,i_{k}}$ for $i_{j}\in[n_{j}]:=\{1,\ldots,n_{j}\}$, where $k$ is called the order of the tensor. For vectors $a^{(1)}\in\mathbb{R}^{n_{1}},\,a^{(2)}\in\mathbb{R}^{n_{2}},\ldots,\,a^{(k)}\in\mathbb{R}^{n_{k}}$, the associated rank-1 tensor is denoted $T=a^{(1)}\otimes a^{(2)}\otimes\cdots\otimes a^{(k)}$ and has entries $T_{i_{1},i_{2},\ldots,i_{k}}=a^{(1)}_{i_{1}}a^{(2)}_{i_{2}}\cdots a^{(k)}_{i_{k}}$. Also, $a^{\otimes k}$ is shorthand for $a\otimes a\otimes\cdots\otimes a$ ($k$ times). While various notions of tensor rank exist, we consider the CP (canonical polyadic) decomposition, that is, we define the rank of a tensor $T$ to be the minimum $r$ such that $T$ can be expressed as the sum of $r$ rank-1 tensors. In the matrix case $k=2$, this coincides with the usual notion of matrix rank.
@@
Three natural and interrelated questions of interest arise when studying decompositions for $T$.
  • Uniqueness Theorem: A uniqueness theorem specifies an explicit list of conditions (ideally, efficiently checkable) on the collection of components $\{a^{(i,\ell)}\,:\,i\in[k],\,\ell\in[r]\}$ that guarantees the uniqueness of the decomposition (1) up to natural ambiguities in specifying the components (see Definition 2.2). E.g., note that we cannot hope to resolve the norm of each component $a^{(i,\ell)}$, since $a\otimes b\otimes c=(2a)\otimes(b/2)\otimes c$.
@@
  • Decomposition: Given a tensor $T$ of the form (1) generated according to some model, the goal is to recover the individual rank-1 terms $a^{(1,\ell)}\otimes\cdots\otimes a^{(k,\ell)}$ up to re-ordering (i.e., the $r$ terms can be listed in any order) and rescalings (again, see Definition 2.2).
Tensors pose a number of computational challenges because basic operations such as computing the rank or finding the best rank-1 approximation are NP-hard for tensors of order 3 and up [Hås90, HL13, Swe18]. As a result, we do not hope for an efficient (i.e., polynomial-time) tensor decomposition algorithm that works on all possible inputs, and it will be important to impose some non-degeneracy assumptions on the input tensor. It is typical to assume the components of $T$ satisfy:
@@
  • Smoothed analysis model: Here, the components are obtained by starting with an arbitrary collection of components and then perturbing them by a small amount of random noise. Additional noise may also be added to the observed tensor. While results in the generic setting tend to assume exact access to the input tensor and exact computation of linear-algebraic primitives, results in the smoothed setting tend to provide a more detailed analysis of the algorithm’s numerical stability and robustness to noise.
@@
Prior work.
@@
First we give an overview of prior works on tensor decomposition, focusing mainly on tensors of order $k=3,4$ and, for ease of exposition, restricting our attention to square tensors, i.e., those of format $n\times n\times\cdots\times n$ with $n$ large. A first question is that of identifiability: when are minimum-rank tensor decompositions unique? A simple argument shows that an order-$k$ tensor of format $n\times\cdots\times n$ cannot have rank larger than $n^{k-1}$ (for instance, an $n\times n\times n$ tensor is comprised of $n$ matrix “slices” and each is $n\times n$ with rank $\leq n$). This means if $T$ is generated according to (1) with $r>n^{k-1}$, the planted decomposition will not be the minimum-rank one, so we do not hope to find it. On the other hand, if $r\leq cn^{k-1}$ for a constant $c=c(k)>0$ and the components are generically chosen, then the planted decomposition will be the unique rank-$r$ decomposition (for $k\geq 3)$ [BCO14]; so in this regime we might hope to solve decomposition (and related problems) in principle, but this does not guarantee an efficient algorithm.
@@
Known polynomial-time algorithms require $r$ to be significantly below the identifiability threshold. For order-3 tensors, a classical method called simultaneous diagonalization or Jennrich’s algorithm can decompose generic tensors of rank $r\leq n$ (the “undercomplete” case) [Har70, LRA93] (see [Moi18]). Furthermore, the analysis comes with a uniqueness theorem: if $T=\sum_{\ell=1}^{r}a^{(\ell)}\otimes b^{(\ell)}\otimes c^{(\ell)}$ where $\{a^{(\ell)}\}$ and $\{b^{(\ell)}\}$ are each linearly independent are $\{c^{(\ell)}\}$ are pairwise linearly independent, then the decomposition is unique and the algorithm will find it. Another well-known uniqueness theorem due to Kruskal [Kru77] has the advantage of tolerating higher rank $r\leq 3n/2$ but the disadvantage that the conditions on the components are not efficiently checkable and no algorithm is provided to find the decomposition; see also [LP23] and references within for various extensions of Kruskal’s theorem.
For order-4 tensors, the FOOBI algorithm and variants [Car91, DL06, DLCC07, MSS16, HSS19, JLV23] can decompose generic tensors of rank $r\leq cn^{2}$ for a constant $c>0$. See [JLV23] for a recent account, which generalizes the setting and corrects an error in the original analysis of [DL06]; this result also certifies uniqueness of the decomposition, providing an efficiently-checkable uniqueness theorem. Noise-robustness and numerical precision of (variants of) the above algorithms have been studied (for instance, in the smoothed analysis model) for order-3 [BCMV14b, GVX14, KS23] and order-4 [MSS16, HSS19] tensors.
@@
If the components are taken to be random rather than generic, striking improvements over the above are possible for order-3 tensors, with poly-time algorithms known to reach $r\approx n^{3/2}$ (up to logarithmic factors in $n$[GM15, MSS16, HSSS16, KP20, DdL+22]. It appears likely that this threshold — and more generally, $n^{k/2}$ for order-$k$ tensors — is the fundamental limit for efficient algorithms in the random model, as evidenced by lower bounds against low-degree polynomials [Wei23]. Note that for order-4 tensors, the algorithms for the generic model match the lower bounds for the (easier) random model, meaning we (likely) cannot expect further improvements in the generic setting, and the generic setting is no harder than the random setting. In contrast, known algorithms for order-3 tensors display a gap between the generic ($r\approx n$) and random ($r\approx n^{3/2}$) settings.
Generic overcomplete order-3 tensors.
@@
The above discussion suggests the possibility of finding improved algorithms for generic $n\times n\times n$ tensors. For some time, the question of finding algorithms (or an efficient uniqueness theorem) in the “overcomplete” regime $r>n$ remained a notable open problem (highlighted at COLT 2014 [BCMV14a] and also mentioned in [MSS16, Moi18]), and recently some progress has been made. The work of [CR22] gave an algorithm for decomposing slightly overcomplete tensors, namely $r=n+\ell$ where the runtime is polynomial in $n$ but exponential in $\ell$. More recently, Koiran [Koi24a] gave both a decomposition algorithm and an efficient uniqueness theorem for $r\leq 4n/3$; this algorithm only uses 4 “slices” of the tensor, and works more generally for tensors of format $n\times n\times p$ for $p\geq 4$.
The current work aims to push these approaches to the limit, improving the condition on $r$ as much as possible and also investigating whether there are fundamental limitations for efficient algorithms. Even a modest improvement on the value of $r$ has potential consequences for downstream applications in learning. For example, consider the heterogeneous multi-reference alignment [PWB+19] problem. Here the goal is to learn a collection of $K$ “signals” ($n$-dimensional vectors), given a mixture of noisy copies of these signals that have been circularly shifted by unknown offsets. When applying the tensor decomposition approach of [PWB+19] using the third moment tensor, the rank of the $n\times n\times n$ tensor that needs to be decomposed is $Kn$. This would be interesting even for $K=2$. Since no such decomposition algorithms are known (except for $K=1$), [PWB+19] instead uses the order-5 moment tensor, which is more expensive in terms of runtime and sample complexity.
Even more tantalizing is the prospect of potentially finding significantly better algorithms for generic tensors, say, matching the threshold $r\approx n^{3/2}$ that appears for random tensors. On one hand, this appears plausible because order-4 tensors have no gap between the generic and random settings, and it would perhaps be unexpected to have such a qualitative difference between odd-order and even-order tensors. For instance, order-$k$ tensor decomposition may appear superficially similar to refutation of $k$-XOR equations — which can be viewed as a tensor-valued problem — where we know the odd- and even-order cases both have the threshold $n^{k/2}$ for both random and smoothed settings, even though the odd-$k$ case is much more technically challenging (see [GKM22]). The same thresholds also arise for the problem of completing low-rank tensors, with similar complexities in the odd-$k$ case [BM16, KP20, LM20]. On the other hand, tensor decomposition appears related to the problem of proving lower bounds on the rank of explicit tensors, which is connected to arithmetic complexity theory (where “explicit” has a precise meaning) [AFT11, Raz13, LO15, Lan15]. Known barriers suggest that it may be hard to go beyond $r\approx n$ for order-3 tensors [EGOW18, GMOW19].
Our contributions: upper and lower bounds.
@@
Our starting point is the work of Garg, Kayal and Saha [GKS20] who showed a surprising connection between algebraic complexity theory and learning. In particular they showed how a powerful lower bound method, called the method of shifted partial derivatives [CKW11], could be leveraged to give algorithms for decomposing power sums of polynomials, which is a close relative of tensor decomposition. Specifically we will be interested in Koszul–Young flattenings [LO13] (see also [Lan11]) which are the main ingredient in recent lower bounds on the border rank of the matrix multiplication tensor [LO15]. Their algorithmic applications were studied in the PhD thesis of Persu [Per18] which gave an efficient algorithm for rank detection (but not decomposition) when $r\leq 3n/2$.
@@
Our algorithmic results (“upper bounds”) apply to order-3 tensors of format $n_{1}\times n_{2}\times n_{3}$, where (without loss of generality) $n_{1}\leq n_{2}\leq n_{3}$. To simplify the discussion, suppose further that we are in an asymptotic regime where $n_{1}\to\infty$ and $n_{3}/n_{2}=O(1)$. In the setting of generic components, we give algorithms for both the rank detection and decomposition tasks, as well as an efficient uniqueness theorem. These results tolerate rank as large as $r\leq(1-\epsilon)(n_{2}+n_{3})$ for an arbitrary $\epsilon>0$. For any fixed $\epsilon$, the runtime is polynomial in $n_{3}$. For comparison, the previous state-of-the-art results for rank detection [Per18] and decomposition [Koi24a] assume $n_{2}=n_{3}=n$ and require $r\leq 3n/2$ and $r\leq 4n/3$ respectively. Also, our efficient uniqueness theorem is the first uniqueness theorem (efficient or not) to surpass Kruskal’s classical bound $r\leq 3n/2$ [Kru77].
@@
In addition to the Koszul–Young flattenings, a key ingredient in our approach will be the algorithm of [JLV23] (see also the precursors [Car91, DL06, DLCC07, BVD+18]) for finding rank-1 matrices in a linear subspace. In contrast, the decomposition algorithm of [Koi24a] takes a different approach, based on connections with the problem of computing a commuting extension. We can exploit this connection in the other direction: Appendix B shows how our decomposition algorithm implies an algorithm for commuting extensions that is (in some sense) nearly optimal.
The blueprint for our algorithm is to solve rank detection by “flattening” the tensor to a particular matrix whose rank can be related to the original tensor rank. We also investigate limitations of this approach (“lower bounds”), building on [EGOW18]. We give a various results for successively broader classes of flattenings. First we identify a subclass of linear flattenings that provably cannot surpass rank $n_{2}+n_{3}$, making our method essentially optimal within this class. We also define a much more general class of degree-$d$ polynomial flattenings and prove that for $n\times n\times n$ tensors, these cannot surpass rank $Cn$ for a constant $C=C(d)$. This extends [EGOW18], which handles the case $d=1$. We note that this style of lower bound is distinct from the low-degree polynomial framework used by [Wei23] for the case of random components, as that approach only applies to fully-Bayesian random models.
@@
Our lower bounds formalize the sense in which reaching $r\gg n$ would require a rather different approach, and might be fundamentally out of reach for efficient algorithms. Of course, there are other approaches that are not ruled out by our lower bounds. For instance, there is the commuting extension approach of [Koi24a], but curiously this appears to have a natural breaking point at $r=2n$ (see Section 2 of [Koi24b]), matching our algorithm. We note that there is an explicit tensor with a rank lower bound of roughly $3n$ [AFT11], but this does not appear to suggest a decomposition algorithm.
@@
We leave for future work the question of obtaining a detailed analysis of the stability and robustness of our algorithm, e.g., in the style of [BCMV14b, GVX14]. Certainly our description and analysis of the algorithm in exact terms is a necessary step toward this goal. Another seemingly necessary step would be to obtain a robust analysis of the [JLV23] algorithm (which we use as a subroutine), but currently this only exists for the “non-planted” variant of the problem [BESV24].

### v2

@@ 1 Introduction
A tensor is a multi-way array $T\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{k}}$ with entries denoted $T_{i_{1},i_{2},\ldots,i_{k}}$ for $i_{j}\in[n_{j}]:=\{1,\ldots,n_{j}\}$, where $k$ is called the order of the tensor. For vectors $a^{(1)}\in\mathbb{R}^{n_{1}},\,a^{(2)}\in\mathbb{R}^{n_{2}},\ldots,\,a^{(k)}\in\mathbb{R}^{n_{k}}$, the associated rank-1 tensor is denoted $T=a^{(1)}\otimes a^{(2)}\otimes\cdots\otimes a^{(k)}$ and has entries $T_{i_{1},i_{2},\ldots,i_{k}}=a^{(1)}_{i_{1}}a^{(2)}_{i_{2}}\cdots a^{(k)}_{i_{k}}$. Also, $a^{\otimes k}$ is shorthand for $a\otimes a\otimes\cdots\otimes a$ ($k$ times). While various notions of tensor rank exist, we consider the CP (canonical polyadic) decomposition, that is, the rank of a tensor $T$ is the minimum $r$ such that $T$ can be expressed as the sum of $r$ rank-1 tensors. In the matrix case $k=2$, this coincides with the usual notion of matrix rank.
@@
Three natural and interrelated questions of interest arise when studying decompositions for $T$:
  • Uniqueness Theorem: A uniqueness theorem specifies an explicit list of conditions (ideally, efficiently checkable) on the collection of components $\{a^{(i,\ell)}\,:\,i\in[k],\,\ell\in[r]\}$ that guarantees the uniqueness of the decomposition (1) up to natural ambiguities in ordering and scaling (see Definition 2.2). E.g., note that we cannot hope to resolve the norm of each component $a^{(i,\ell)}$, since $a\otimes b\otimes c=(2a)\otimes(b/2)\otimes c$.
@@
  • Decomposition: Given a tensor $T$ of the form (1) generated according to some model, the goal is to recover the components $a^{(i,\ell)}$, again up to ordering and scaling.
Tensors pose a number of computational challenges because basic operations such as computing the rank or finding the best rank-1 approximation are NP-hard for tensors of order 3 and up [Hås90, HL13, Swe18]. As a result, we do not hope for an efficient (i.e., polynomial-time) tensor decomposition algorithm that works on all possible inputs, and it will be important to impose some non-degeneracy assumptions on the input tensor. Some common assumptions include:
@@
  • Smoothed analysis model: Here, the components are obtained by starting with an arbitrary collection of components and then perturbing them by a small amount of random noise. Additional noise may also be added to the observed tensor. While results in the generic setting tend to assume exact access to the input tensor and exact computation of linear-algebraic primitives, results in the smoothed setting also need to address the algorithm’s numerical stability and robustness to noise.
@@
1.1 Prior Work
Generic Case.
@@ 1.1 Prior Work
First we give an overview of prior work on third-order tensor decomposition, focusing on square tensors, i.e. those of format $n\times n\times n$, with $n$ large. A first question is that of identifiability: when are minimum-rank tensor decompositions unique? A tensor of format $n\times n\times n$ cannot have rank larger than $n^{2}$, because it is comprised of $n$ matrix “slices” and each is $n\times n$ with rank $\leq n$. This means if $T$ is generated according to (1) with $r>n^{2}$, the planted decomposition will not be the minimum-rank one, so we do not hope to find it. On the other hand, if $r\leq cn^{2}$ for a particular constant $c\in(0,1)$ and the components are generically chosen, then the planted decomposition will be the unique rank-$r$ decomposition [BCO14]. See also [COV14] for identifiability results for tensors of specific sizes.
@@
But what about computationally efficient algorithms? Known polynomial-time algorithms require $r$ to be significantly below the identifiability threshold. For order-3 tensors, a classical method called simultaneous diagonalization or Jennrich’s algorithm can decompose generic tensors of rank $r\leq n$ (the “undercomplete” case) [Har70, LRA93] (see [Moi18]). Furthermore, the analysis comes with a uniqueness theorem: if $T=\sum_{\ell=1}^{r}a^{(\ell)}\otimes b^{(\ell)}\otimes c^{(\ell)}$ where $\{a^{(\ell)}\}$ and $\{b^{(\ell)}\}$ are each linearly independent and $\{c^{(\ell)}\}$ are pairwise linearly independent, then the decomposition is unique and the algorithm will find it. Another well-known uniqueness theorem due to Kruskal [Kru77] has the advantage of being applicable even for rank up to $r\leq 3n/2$ but the disadvantage that the conditions on the components are not efficiently checkable and no algorithm is known for finding the decomposition in polynomial time; see also [LP23, Gub24] and references within for various extensions of Kruskal’s theorem.
Random Case.
If the components are taken to be random rather than generic, there are efficient algorithms that succeed for much larger values of $r$. Known algorithms reach $r\approx n^{3/2}$ (up to logarithmic factors in nn[GM15, MSS16, HSSS16, KP20, DdL+22]. However these algorithms exploit the randomness of the components in an essential way: They rely on the observation that the maxima of $\langle T,x\otimes y\otimes z\rangle$ over unit vectors $x$, $y$ and $z$ occur near the unknown components. Thus many algorithms for decomposing random third-order tensors start from a semidefinite programming relaxation for the above polynomial optimization problem and show how to round its solutions to recover the components. However this strategy quickly breaks down even in the smoothed model because the magnitude of the perturbations can be smaller than the magnitude of the components in the base tensor by a large polynomial factor.
@@
It appears likely that this threshold — and more generally, $n^{k/2}$ for order-$k$ tensors — is the fundamental limit for efficient algorithms in the random model, as evidenced by lower bounds against low-degree polynomials [Wei23]. We remark that for fourth-order tensors (discussed in Section 1.3), the algorithms for the generic model match the lower bounds for the (easier) random model, meaning we (likely) cannot expect further improvements in the generic setting, and the generic setting is no harder than the random setting. In contrast, known algorithms for third-order tensors have a gap between the generic ($r\approx n$) and random ($r\approx n^{3/2}$) settings.
Generic third-order overcomplete tensor decomposition.
Our main focus will be on algorithms for decomposing overcomplete third-order tensors in the generic case. We remark that even a modest improvement on the value of $r$ has potential consequences for downstream applications in learning, since it often imposes an upper bound on the complexity of models we can work with (e.g., the number of components in a mixture model). And the assumption that the components are random often leads to much more stringent assumptions on the model as well, at least compared to the generic case.
@@
For some time, the question of finding any algorithm (or an efficient uniqueness theorem) in the “overcomplete” regime $r>n$ remained a notable open problem (highlighted at COLT 2014 [BCMV14a] and also mentioned in [MSS16, Moi18]), and recently some progress has been made. The work of [CR22] gave an algorithm for decomposing slightly overcomplete tensors, namely $r=n+\ell$ where the runtime is polynomial in $n$ but exponential in $\ell$. More recently, Koiran [Koi24a] gave both a decomposition algorithm and an efficient uniqueness theorem for $r\leq 4n/3$; this algorithm only uses 4 “slices” of the tensor, and works for tensors of format $n\times n\times p$ for $p\geq 4$. Our work aims to push these approaches to the limit, improving the condition on $r$ as much as possible and also investigating whether there are fundamental limitations for efficient algorithms.
1.2 Our Contributions: Upper and Lower Bounds.
@@
Our starting point is the work of Garg, Kayal and Saha [GKS20] who showed a surprising connection between algebraic complexity theory and learning. In particular they showed how a powerful lower bound method, called the method of shifted partial derivatives [CKW11], could be leveraged to give algorithms for decomposing polynomials into power sums, which is a close relative of tensor decomposition. We will similarly build our algorithms out of certain algebraic gadgets, which we formalize below:
Definition 1.1.
A rank detection gadget (a.k.a. “flattening”) $M$ is an efficiently computable function that maps third-order tensors $T\in\mathbb{R}^{n\times n\times n}$ to matrices $M(T)$. We say that $M$ solves the generic rank detection problem up to rank $r$ if there is some function $f$ such that:
  1. 1.
    For all third-order tensors $T$, $f(\mathrm{rank}(M(T))\leq\mathrm{rank}(T)$.
  2. 2.
    Moreover for generic tensors $T$ up to rank $r$, $f(\mathrm{rank}(M(T))=\mathrm{rank}(T)$.
This definition is inspired by the definition of what it means to refute a random CSP. In particular we want that $f(\mathrm{rank}(M(T))$ should always yield a valid lower bound on the rank of $T$, but we only require this lower bound to be accurate on generic tensors up to some given rank. The main question is: How can we design rank detection gadgets that work for the largest range of ranks?
A simple rank detection gadget is to just take the first “slice” through the tensor. It is easy to check that its rank is always a lower bound on the rank of $T$. But since $M(T)$ is an $n\times n$ matrix, this gadget can only plausibly work for generic tensors up to rank $n$. Devising rank detection gadgets that work beyond $n$, i.e. in the overcomplete regime, turns out to be rather non-trivial. The first rank detection gadget that plausibly works beyond $n$ was given by Strassen [Str83]. The key property is that if $T$ is $n\times n\times n$ then $M(T)$ is a $3n\times 3n$ matrix and moreover the mapping is linear and sends a rank-one tensor $T$ to a rank-two matrix $M(T)$. Thus Strassen’s gadget plausibly works up to rank $r\leq 3n/2$. But proving this requires showing that the rank two matrices that arise from each component do not generically interfere with each other. Strassen’s gadget was analyzed in the PhD thesis of Persu [Per18] who showed it solves rank detection (but not decomposition) when $r\leq 3n/2$. It turns out this is an example of an even more general strategy called Koszul–Young flattenings [LO13] (see also [Lan11]), which are the best known rank detection gadgets. These are parameterized by a positive integer $p$ and map an $n\times n\times n$ tensor $T$ to a $\binom{2p+1}{p}n\times\binom{2p+1}{p}n$ matrix $M(T)$. They are again linear, and map a rank-one tensor $T$ to a rank-$\binom{2p}{p}$ matrix $M(T)$, so they plausibly work up to $r=(2-\epsilon)n$ for any $\epsilon>0$.
Our main algorithmic results are tensor decomposition and efficient uniqueness theorems that reach the same limit as the Koszul–Young flattenings. Below we give an informal and simplified statement of the main algorithm results (“upper bounds”), with the full details presented in Section 2.2.5.
Theorem (Upper Bounds, Informal).
@@
Consider tensors of format $n_{1}\times n_{2}\times n_{3}$ where (without loss of generality) $n_{1}\leq n_{2}\leq n_{3}$, and further assume an asymptotic regime where $n_{1}\to\infty$ and $n_{3}/n_{2}=:\alpha=O(1)$. In the setting of generic components, we give algorithms for both the rank detection and decomposition tasks, as well as an efficient uniqueness theorem. These results tolerate rank as large as $r\leq(1-\epsilon)(n_{2}+n_{3})$ for an arbitrary constant $\epsilon>0$. For any fixed $\epsilon$, the runtime is polynomial in $n_{3}$. More precisely, the runtime is polynomial in $n_{3}\cdot 2^{\alpha/\epsilon}$.
For comparison, the previous state-of-the-art results for rank detection [Per18] and decomposition [Koi24a] assume $n_{2}=n_{3}=n$ and require $r\leq 3n/2$ and $r\leq 4n/3$ respectively. Also, our efficient uniqueness theorem is the first uniqueness theorem (efficient or not) to surpass Kruskal’s classical bound $r\leq 3n/2$ [Kru77].
@@
In addition to the Koszul–Young flattenings, a key ingredient in our approach will be the algorithm of [JLV23] (see also the precursors [Car91, DL06, DLCC07, BVD+18]) for finding rank-1 matrices in a linear subspace. In contrast to our approach based on rank detection gadgets, the decomposition algorithm of [Koi24a] takes a different approach, based on connections with the problem of computing a commuting extension. We can exploit this connection in the other direction: Appendix B shows how our decomposition algorithm implies an algorithm for commuting extensions that is (in some sense) nearly optimal.
We also investigate limitations of our approach (“lower bounds”) by showing that certain natural strategies cannot even solve the easier rank detection problem. We give a few different results for successively broader classes of flattenings. First we consider linear flattenings of the form $M(a\otimes b\otimes c)=A(a)\otimes(bc^{\top})$ for a matrix $A(a)$ whose entries are homogeneous degree-1 polynomials in the entries of $a$. We designate such flattenings as Koszul–Young type because the Koszul–Young flattenings take this form. We also consider the substantially more general degree-$d$ polynomial flattenings where each entry of $M(T)$ is a homogeneous degree-$d$ polynomial in the entries of $T$. In both cases, we do not quite rule out rank detection in the sense of Definition 1.1, but we rule out a natural sufficient condition called rank additivity. In the case of linear flattenings, rank additivity corresponds to Definition 1.1 with $f(x)=mx$ for some $m\in\mathbb{Z}_{\geq 1}$. Below we give an informal and simplified statement of our lower bounds, with the full details presented in Section 2.3.
Theorem (Lower Bounds, Informal).
Consider tensors of format $n_{1}\times n_{2}\times n_{3}$ where $n_{1}\leq n_{2}\leq n_{3}$.
  • Flattenings of Koszul–Young type can only exhibit rank additivity when $r\leq n_{2}+n_{3}$.
  • Degree-$d$ homogeneous polynomial flattenings can only exhibit rank additivity when $r\leq Cn_{3}$ for a constant $C=C(d)$.
The bound $n_{2}+n_{3}$ in the first result is precisely the threshold achieved by our algorithm, so the Koszul–Young flattenings are optimal among flattenings of this style. The proofs extend [EGOW18], which covers the case $d=1$ of the second result. We note that this style of lower bound is distinct from the low-degree polynomial framework used by [Wei23] for the case of random components, as that approach only applies to fully-Bayesian random models.
@@
Our lower bounds formalize the sense in which reaching $r\gg n$ would require a rather different approach, and might be fundamentally out of reach for efficient algorithms. Of course, there are other potential approaches that are not ruled out by our lower bounds. For instance, we have not ruled out a flattening whose rank tracks that of the original tensor in a non-additive way. Beyond flattenings, there is the commuting extension approach of [Koi24a], but curiously this appears to have a natural breaking point at $r=2n$ (see Section 2 of [Koi24b]), matching our algorithm. We note that there is an explicit tensor with a rank lower bound of roughly $3n$ [AFT11], but this does not appear to suggest a decomposition algorithm.
@@
We leave for future work the question of obtaining a detailed analysis of the stability and robustness of our algorithm, e.g., in the style of [BCMV14b, GVX14]. Certainly our description and analysis of the algorithm in exact terms is a necessary step toward this goal. Another seemingly necessary step would be to obtain a robust analysis of the [JLV23] algorithm (which we use as a subroutine), but currently this only exists for the “non-planted” variant of the problem [BESV24] (where the goal is to refute the existence of a rank-1 matrix in the given subspace).
1.3 Further Discussion
We remark that there are fourth-order tensor decomposition algorithms that work in the generic and highly overcomplete setting. In particular the FOOBI algorithm and its variants [Car91, DL06, DLCC07, MSS16, HSS19, JLV23] can decompose generic tensors of rank $r\leq cn^{2}$ for a constant $c>0$. See [JLV23] for a recent account, which generalizes the setting and corrects an error in the original analysis of [DL06]; this result also certifies uniqueness of the decomposition, providing an efficiently-checkable uniqueness theorem. However from the vantage point of Definition 1.1 an associated rank detection gadget for fourth-order tensors is easy to construct: We can rearrange the $n\times n\times n\times n$ tensor into an $n^{2}\times n^{2}$ matrix. This works as a rank detecting gadget for rank $r\leq n^{2}$. Also noise-robustness and numerical precision of (variants of) the above algorithms have been studied (for instance, in the smoothed analysis model) for order-3 [BCMV14b, GVX14, KS23] and order-4 [MSS16, HSS19] tensors.